# Laminator: Verifiable ML Property Cards using Hardware-assisted Attestations

### v1

@@ VII Related Work
ML for Auditing. Several ML based techniques can be used for auditing ML models. For instance, membership inference attacks have been repurposed to be used for auditing by checking if data, requested to be removed as part of right to be forgotten, belongs to training data [58, 42]. However, such membership inference attacks have been shown to be unreliable and thereby lack effectiveness [53, 35]. Similarly, distribution inference attacks can be repurposed to be used for inference-based property attestation [11, 30]. However, these approaches lack effectiveness as well [11]. Further, Jia et al. [28] propose proof of training to identify whether a model was trained on a specific training dataset. However, several subsequent works have shown that these approaches can be easily evaded and are not effective [15, 69]. Hence, current ML based approaches for auditing are not reliable and robust.
Certified ML. Several prior works provide a “certificate” which is basically a bound on the accuracy in presence of adversarial examples (see a survey by Li et al. [39]). Further, implementation of differential privacy could have bugs due to which the actual privacy bounds are not the same as the claimed bounds. Several works have been proposed to audit differentially private models [60, 27, 48]. Finally, bounds to guarantee group and individual fairness can also be generated [29, 33, 65]. These bounds on different ML properties can be generated using randomized smoothing over the inputs or model parameters, or using formal theory. However, the threat model for these approaches are different from ours: these works implicitly assume that the prover who generates the proofs is honest while we assume the prover can be malicious. We instead rely on hardware as the root of trust.
@@
Cryptographic Primitives for Auditing. ZKPs have been extensively used for verifiable inference to ensure that the model’s output has not been tampered [40, 16, 67, 24, 70, 19]. ZKPs for training are still in early stages and have been used for proof of training for different models [62, 57, 17, 5, 13]. However, this is not very efficient as proofs have to be generated for each of the training epoch which limits it to simple ML models [57, 17]. Extending these approaches to neural networks is an open problem as it incurs a significant overhead [62, 5]. Similarly, MPC for proof of training requires to train model for each attestation thereby incuring a high overhead [11]. Further, cryptographic primitives can verify properties such as fairness [57, 34, 50, 55, 68] and privacy [56]. Srivastava et al. [59] propose a verifiable training scheme aims to ensure determinism, mitigating hardware discrepancies and improving performance for auditing [17]. Finally, MPC has been used to verify entire training and inference pipeline but require multiple interactions [41].
Confidential Computing for ML. TEEs have been extensively explored for both training [45] and inference of ML models [46] (see survey by Mo et al. [47]). Prior works have relied on different optimizations including model partitioning to get ML models to execute within TEEs. Further, TEEs can be combined with GPUs for effiicent execution by using additional cryptographic operations to check for the integrity of the data transferred between TEE and GPU [64]. However, most these worst focus on ensuring confidentiality of the model and inputs during inference or training datasets. None of them use TEEs for integrity to furnish attestations.
VIII Discussion and Summary
Completeness. We consider three types of ML property cards and different attestations for each of them as illustrative examples in Section IV. However, we do not claim them to be complete. There could be other attestations and ML property cards but Laminator can be extended to account for them (discussed in Section V-A). For instance, ecosystem cards which considers a combination of multiple models trained on different datasets [6]. Hence, they comprise of multiple datasheets and model cards. However, Laminator can make them verifiable using verifiable model cards and datasheets.
Attacks on Intel SGX. Several prior works have shown that Intel SGX is susceptible to a wide range of side channel attacks by which an adversary can violate confidentiality [66]. However, we assume that model provider already has access to the training dataset and model details. There is no requirement for confidentiality but they use Laminator to show that their training and inference process is as expected. Hence, leakage via side channels is not a concern in our setting. The exception to this is that leakage of attestation keys will allow violation of the integrity of the attestation [7, 66]. Hence, Laminator relies on the ability of the platform developer to ensure the integrity of the attestation, using hardware and software mechanisms to prevent subversion e.g. by using side-channel or other attacks to extract attestation keys. However, this is an orthogonal problem to our own and the subject of active research and development effort on the part of TEE manufacturers.
Improving Efficiency with GPUs. We chose TEEs as the new variants have larger memory and rely on AMX extensions which can allow for faster inference and training. To improve efficiency, similar to prior works [64, 71], we can outsource the execution of some of the model layers to a GPU. While these works focus on confidentiality, we have to check for integrity of the computation done on the GPU. Further, using the improvements in hardware to combine GPUs with TEE (e.g., Nvidia’s H100 has a new TEE with a GPU), $\mathcal{P}$ can adapt Laminator to the new hardware to run both training and inference securely inside TEEs to improve efficiency [49]. This can allow to evaluate larger models including large language models.
Summary. ML property cards are widely adapted by companies such as Huggingface. There is a need to ensure the information claimed in these cards are correct creating the need for verifiable ML property cards. We propose Laminator, the first framework to generate verifiable ML property cards using hardware-assisted ML property attestations. We evaluate Laminator and show that it satisfies efficiency, scalability and versatility.
IX Acknowledgements
@@
This work is supported in part by the Research Council of Finland (decision 339514), Intel Labs, and the Government of Ontario. Views expressed in the paper are those of the authors and do not reflect the position of the funding agencies.

### v3

@@ 7. Related Work
ML-based Auditing. Several ML-based techniques, such as membership inference attacks, are used for auditing models by checking if data requested for removal under the right to be forgotten belongs to the training set (Song and Shmatikov, 2019; Miao et al., 2021). Distribution inference attacks can also be used for inference-based property attestation (Duddu et al., 2024; Juarez et al., 2022). Both membership inference and distribution inference are unreliable and ineffective for auditing (Rezaei and Liu, 2021; Kong et al., 2022; Duddu et al., 2024). Jia et al. (Jia et al., 2021) propose proof of training to verify if a model was trained on a specific dataset. However, it can be easily evaded (Fang et al., 2023; Zhang et al., 2022). Thus, ML-based techniques are unreliable for auditing.
Certified ML. Several works provide “certificates” as accuracy bounds in the presence of adversarial examples (e.g., Li et al. (Li and Xothers, 2023)). Differential privacy implementations may have bugs, leading to mismatched privacy bounds, prompting audits of such models (Steinke et al., 2024; Jagielski et al., 2020; Nasr et al., 2023). Additionally, bounds for group and individual fairness can be generated (Jin et al., 2022; Khedr and Shoukry, 2023; Urban et al., 2020). These bounds are generated using randomized smoothing or formal theory. This is an orthogonal problem as attestation attestation guarantees the computations done, not their correctness. For instance, attestation of differential privacy training only confirms that differential privacy was used, not whether its implementation or guarantee is correct.
@@
Cryptographic Primitives for Attestation. ZKPs have been widely used for verifiable inference to ensure model outputs are untampered (Liu et al., 2021; Feng et al., 2021; Weng et al., 2021; Huang et al., 2022; Zhang et al., 2023; Ghodsi et al., 2017). ZKPs for training, though in early stages, have been used for proof of training (Sun and Zhang, 2023; Shamsabadi et al., 2023; Garg et al., 2023; Abbaszadeh et al., 2024; Eisenhofer et al., 2022). However, generating proofs for each training epoch is inefficient, limiting this approach to simple ML models (Shamsabadi et al., 2023; Garg et al., 2023). Extending these approaches to neural networks is an open problem due to significant overhead (Sun and Zhang, 2023; Abbaszadeh et al., 2024). Similarly, MPC for proof of training requires training the model for each attestation, incurring high overhead (Duddu et al., 2024). Cryptographic primitives can also verify properties like fairness (Shamsabadi et al., 2023; Kilbertus et al., 2018; Park et al., 2022; Segal et al., 2021; Yadav et al., 2024) and privacy (Shamsabadi et al., 2024). Srivastava et al. (Srivastava et al., 2024) propose a verifiable training scheme to ensure determinism, reduce hardware discrepancies, and improve auditing performance (Garg et al., 2023). MPC has also been used to verify the entire training and inference pipeline, but requires multiple interactions (Lycklama et al., 2024).
Confidential Computing for ML. TEEs have been widely explored for both training (Mo et al., 2021) and inference of ML models (Mo et al., 2020) (see survey by Mo et al. (Mo et al., 2022)), often using optimizations like model partitioning. TEEs can also be combined with GPUs for efficient execution, adding cryptographic checks for data integrity between TEE and GPU (Tramer and Boneh, 2019). Prior TEE-based approaches focus on data confidentiality (e.g., private inference) (Mo et al., 2022), but ML verifiability requires integrity guarantees. We are not aware of any prior ML verifiability approaches using TEEs.
8. Discussion and Summary
Completeness. We illustrate three types of ML property cards and four type of model attestations in Section 4, but do not claim completeness. There could be other attestations types and ML property cards but Laminator can be extended to account for them by modifying the “measurer” script in the prover (Section 5.1). Any property that can be expressed in Python code can be attested using Laminator. For instance, Laminator can be extended to include other attestation types, such as ecosystem cards, a combination of multiple models trained on different datasets, using verifiable model cards and datasheets (Bommasani et al., 2023). Also, we can attest various steps in DPSGD training within our framework to demonstrate to a verifier that the prover has used the algorithm. Model explanations (during inference) will be handled like Input-Model-Output attestation. While we focus on static properties for various ML operations, there are other runtime properties such as environmental impact of running ML models inside GPUs (e.g., carbon emission). We leave the design of attestations for such runtime properties for future.
Security Analysis and Attacks on TEEs. Laminator does not introduce any cryptographic protocols. Its security guarantees are inherited directly from the underlying attestation mechanism. Several prior works have demonstrated attacks on TEEs, particularly side-channel attacks, that can compromise confidentiality (Van Schaik et al., 2024). Attestation does not introduce new privacy risks, as no additional sensitive data is added to the TEE. Laminator relies on the integrity of the underlying TEE attestation mechanism to provide integrity guarantees for property attestation, complementing the confidentiality properties already offered by the TEE. Hence, leakage of model and training data via side channels is not a concern in our setting. Leakage of attestation keys can compromise the integrity of the attestation mechanism (Bulck et al., 2018; Van Schaik et al., 2024), but this is orthogonal to our approach and not affected by Laminator’s design. Mitigating such risks through side-channel-resistant code (Borrello et al., 2021), cache partitioning (Costan et al., 2016), or other techniques (Yu et al., 2020) is an active area of research by TEE manufacturers. Hence, Laminator relies on the platform developer to ensure attestation integrity, using hardware and software mechanisms to prevent subversion, such as side-channel or other attacks to extract attestation keys.
Improving Efficiency with GPUs. Our idea is independent of the specific TEE in which the computation is attested. Most current TEEs are CPU-based, while most ML tasks use GPUs. New TEE-enabled GPUs like Nvidia’s H100 (Nvidia, 2024) enable CPU-based TEEs to accelerate computations, allowing evaluation of larger models, including language models, with comparable performance (Zhu et al., 2024). We plan to extend support to H100 GPUs. CPU extensions like Intel AMX (Intel, 2024) can accelerate ML operations within existing TEEs, enabling faster inference and training than traditional CPU implementations. Where the performance of CPUs are insufficient, hybrid approaches like Slalom (Tramer and Boneh, 2019) and TEESlice (Zhang et al., 2024) enable a CPU-based TEE to offload model layers to a GPU, using cryptography to ensure data confidentiality and output integrity. This speeds up computation compared to CPU-only execution, but the added cryptographic overhead and data transfer between CPU and GPU reduce performance compared to TEE-enabled GPUs.
Summary. ML property cards, widely used by companies like Huggingface, require verification to ensure their correctness. We propose Laminator, the first framework for generating verifiable ML property cards using hardware-assisted attestations. We demonstrate that Laminator is efficient, scalable, and versatile, enabling higher-performance, verifiably trustworthy ML-based services.
{arxiv}
Acknowledgments
@@
This work is supported in part by the Research Council of Finland (Decision 339514), Intel Labs, the Government of Ontario, and the Natural Sciences and Engineering Research Council of Canada (grant number RGPIN-2020-04744). Vasisht is supported by IBM PhD Fellowship, David R. Cheriton Scholarship, and Cybersecurity and Privacy Excellence Graduate Scholarship. The views expressed in the paper are those of the authors and do not reflect the position of the funding agencies. The authors thank Prach Chantasantitam and Asim Waheed for their help running the code.